OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30). 10 0 5 −10

OpenStudy (diyadiya):

Slope of a line passing through two points (x1,y1) & (x2,y2) is \[\Large Slope = \frac{y2-y1}{x2-x1}\]

OpenStudy (diyadiya):

Here ; (x1,y1) = (0,0) (x2,y2)=(3,30)

OpenStudy (diyadiya):

can you tell me what is x1 ,y1 ,x2 & y2 here?

OpenStudy (anonymous):

\[\frac{30-0}{3-0}=10 \]

OpenStudy (anonymous):

y = 10 x